OpenStudy (anonymous):

I need to find the iverse of the function f(x)=9+sqrt3(x-9)

OpenStudy (anonymous):

\[f(x)=9+\sqrt[3]{x}\]

OpenStudy (anonymous):

Not x-9 my bad.

OpenStudy (anonymous):

y=9+sqrt(x)^3 y-9=sqrt(x)^3 (y-9)^3=x

OpenStudy (anonymous):

Oh I need to solve for y though right because you switch the x and y so it looks like \[x=9+\sqrt[3]{y}\] and then solve for y right?